# Revisiting the Orbital Dynamics of the Hot Jupiter WASP–12b with New Transit Times

### v1

@@ 3.1.1 Data Reduction of high-precision TESS data
We retrieved TESS light curves from the NASA funded astronomical data archive repository Mikulski Archive for Space Telescopes (MAST) [^1] All the TESS data used in this paper can be found in MAST (MAST Team, 2021)., using the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) products generated by the Science Processing Operations Center (SPOC; J. M. Jenkins et al. 2016). Compared to standard SAP data, PDCSAP light curves reduce scatter, suppress short-term noise, and correct long-term instrumental trends (J. C. Smith et al. 2012; M. C. Stumpe et al. 2012, 2014; D. A. Caldwell et al. 2020).
@@
Data pre-processing, including extraction, normalization, and detrending, was performed with the JULIET package (N. Espinoza et al. 2019), which employs Bayesian inference through the MultiNest algorithm (F. Feroz et al. 2009, 2019; J. Buchner 2016). We retained only points with a quality flag of zero and converted timestamps to $\mathrm{BJD_{TDB}}$ by adding 2,457,000.
@@
Stellar and instrumental variability was modeled using a Gaussian Process (GP) with a Matérn kernel implemented in CELERITE (N. Espinoza et al. 2019). A normal prior was set for the mean out-of-transit flux, the dilution factor fixed to unity, and wide log-uniform priors adopted for GP hyperparameters (amplitude, timescale, and jitter).
For transit modeling, in-transit points were masked, the GP was fitted to out-of-transit data, and residuals were removed to produce detrended, normalized light curves. Transit windows were then extracted within $\pm 0.1$ days of the predicted mid-transit times. Incomplete or noisy transits were excluded, as missing ingress or egress hampers precise timing (S. C. C. Barros et al. 2013). This procedure yielded GP-corrected light curves, optimized for precise transit timing and exoplanet characterization.
3.2 New Ground-Based Light Curves
@@
For WASP-12b, we obtained one transit on 8 January, 2025 using the 1.3 m Devasthal Fast Optical Telescope (DFOT) at the Devasthal, Nainital campus of ARIES. In addition, three transits were observed on 26 and 27 December, 2024 and 24 February, 2025 with the 0.61 m reflecting telescope, VASISTHA at the IERCOO campus of ICSP, Kolkata. Three further transits were recorded on 2 January, 4 January, and 26 March 2025 with the 0.3 m AG Optical IDK telescope at the Utah Desert Remote Observatory (UDRO) in Beryl, Utah. The exposure times were optimized to compensate for variable weather conditions. The data reduction procedures have been described in detail in Section 3.2.4.
3.2.1 Devasthal Fast Optical Telescope
@@
The observations of one new transit of WASP-12b were conducted using the 1.3-m Devasthal Fast Optical Telescope (DFOT), located at the Aryabhatta Research Institute of Observational Sciences (ARIES), Nainital, India. The DFOT, a modern Ritchey–Chrétien Cassegrain telescope with a 1.3 m aperture, was installed at Devasthal by DFM Engineering Inc., USA, and is operated by ARIES, an autonomous institute under the Department of Science and Technology (DST), Government of India. Equipped with a fast $f/4$ optical system, the telescope is mounted on a fork–equatorial system with single-axis tracking. Its secondary mirror is controlled by a five-axis actuator for precise focusing, while friction drives provide backlash-free motion in right ascension and declination, ensuring a pointing accuracy of $\sim 10$ arcsec (rms). For imaging, a back-illuminated CCD camera with $2048\times 2048$ pixels ($13.5\,\mu\mathrm{m}$ pixel size) and deep thermoelectric cooling down to $80^{\circ}$C was used. All transit observations were carried out using an $R$-band filter.
3.2.2 Ionospheric and Earthquake Research Centre and Optical Observatory
@@
We observed three transits of the hot Jupiter WASP-12b with the 0.61 m VASISTHA reflecting telescope at the Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), established by ICSP at Sitapur, Paschim Medinipur, West Bengal. This telescope, the largest in eastern India, is mounted on an Ascension 200 German equatorial system with high-resolution encoders. All observations were conducted in the Cousins $R$ band to reduce limb darkening, color-dependent extinction, and to allow high-cadence photometry (M. J. Holman et al. 2007). The telescope employs a 24 inch primary mirror with a focal ratio of $f/6.5$ and is equipped with an Atik 460EX Mono CCD camera. The detector has $2749\times 2199$ pixels of size $4.54~\mu\mathrm{m}$, providing an image scale of $0.235~\mathrm{arcsec~pixel}^{-1}$.
3.2.3 Utah Desert Remote Observatory
@@
3.2.4 Data Reduction of New Ground-based Data
@@
Before performing aperture photometry, we preprocessed the raw CCD images to remove instrumental and observational artifacts. We corrected defects such as bad pixels, cosmic ray hits, and pixel to pixel sensitivity variations to ensure reliable photometric measurements. Along with the science frames, we obtain calibration frames in the form of bias, dark, and flat fields to eliminate readout noise, thermal current, and uneven detector response. Using IRAF [^2] The Image Reduction and Analysis Facility (IRAF) is a software package distributed by the National Optical Astronomy Observatory (NOAO), which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation (NSF). (Image Reduction and Analysis Facility), we apply standard routines such as zerocombine, flatcombine, and ccdproc under the imred package of NOAO to perform trimming, bias subtraction, dark correction, and flat–fielding, and we remove cosmic rays with the cosmic rays task in crutil. After calibration, we inspected the images for residual defects and alignment issues, displayed them with the task display, and interactively analyzed them using imexamine to estimate parameters such as the full width at half maximum (FWHM) of stellar point spread functions and the local sky background. These parameters are essential for defining the photometric aperture radius and sky annulus, which directly affect the precision of the extracted light curves in exoplanet transit analysis.
@@
: We performed the aperture photometry using the apphot and daophot packages in IRAF. First, we defined the photometry parameters in datapars, centerpars, fitskypars, and photpars to specify CCD characteristics, centering methods, sky background estimation, and aperture sizes. We prepared a coordinate file for the target star and nearby comparison stars, although stars could also be selected interactively. Aperture photometry was then applied across all science frames using the phot task to extract instrumental magnitudes, with magnitudes computed for the chosen aperture(s). The aperture radii were typically set to 2-3 times the full width at half maximum (FWHM) of the stellar point-spread function (PSF) and optimized to minimize scatter in the out-of-transit (OOT) light curve. We tracked stellar centroids across all frames, typically by cross-correlating with a reference image, to ensure precise aperture alignment during extraction. Photometry was then performed with the daophot package in IRAF on the target star and 2-8 nearby comparison stars of similar brightness and color. The aperture configuration and number of comparison stars were adjusted slightly from night to night according to atmospheric conditions and field orientation, following the approach of I.-G. Jiang et al. (2016).
@@
Once we extracted the instrumental magnitudes, we compiled the results using txdump and related tools to obtain magnitudes, errors, and centroids for the target and comparison stars. For exoplanet transit studies, we performed differential photometry by constructing a reference flux from the summed fluxes of the chosen comparison stars and dividing the target star’s flux by this reference. This produced a time series of relative flux, forming the transit light curve. To correct for atmospheric transparency variations, we normalized the light curves by fitting a linear function to the out-of-transit (OOT) data, yielding an OOT flux level close to unity. We then analyzed the normalized differential light curve for transit signals, where the characteristic dip in brightness corresponds to the planet passing in front of its host star. Finally, we converted all time stamps to the Barycentric Julian Date in the Barycentric Dynamical Time system (BJDTDB) using the publicly available code of J. Eastman et al. (2010). Further modeling and fitting, allowed us to refine the stellar and planetary parameters of the exoplanet, which are described in detail in Section 4.
@@
Object Name Telescope Epoch TDB-based BJD Normalized Flux Normalized Flux Error
WASP-12 1.3 m 5652 2460684.12708776 1.0015969 0.0031616
WASP-12 1.3 m 5652 2460684.12749167 1.005571 0.0031616
WASP-12 1.3 m 5652 2460684.12789558 1.0047181 0.0031616
WASP-12 0.61 m 5640 2460671.05890418 1.0072065 0.0082231
WASP-12 0.61 m 5640 2460671.05989713 1.0388619 0.00822
WASP-12 0.61 m 5640 2460671.06088998 1.029499 0.008217
WASP-12 0.61 m 5641 2460672.13016962 0.9942078 0.0055637
WASP-12 0.61 m 5641 2460672.13116258 1.0037198 0.005564
WASP-12 0.61 m 5641 2460672.13215611 1.0047568 0.0055644
WASP-12 0.61 m 5695 2460731.04994532 1.0013464 0.0076853
WASP-12 0.61 m 5695 2460731.05015155 1.006992 0.0076853
WASP-12 0.61 m 5695 2460731.05033927 1.0045774 0.0076853
WASP-12 0.3 m 5647 2460678.655638 1.0089026 0.0034889
WASP-12 0.3 m 5647 2460678.6564131 1.0023692 0.0034891
WASP-12 0.3 m 5647 2460678.65752766 0.9950521 0.0034893
WASP-12 0.3 m 5648 2460680.88122133 1.0026334 0.0061924
WASP-12 0.3 m 5648 2460680.88228234 0.9922296 0.0061924
WASP-12 0.3 m 5648 2460680.88334356 0.9953348 0.0061924
WASP-12 0.3 m 5723 2460761.64856153 0.9988078 0.0060729
WASP-12 0.3 m 5723 2460761.65071931 0.9930364 0.0060729
WASP-12 0.3 m 5723 2460761.65282303 1.0030208 0.006073
@@
3.3 Ground-based Observations from Public Databases
3.3.1 Exoplanet Transit Database
The ETD, established in September 2008, serves as a collaborative platform where amateur astronomers worldwide contribute transit measurements. As of 29 September, 2025, the database contains 83,053 observations contributed by 1636 observers from different observatories across the globe. In addition to the TESS observations, we incorporated complete transit light curves from the Exoplanet Transit Database (ETD), selecting only those with a data quality index ($DQ<3$) to ensure consistency and reliability. These high-quality, community sourced datasets provided an essential complement to the TESS data, thereby improving the robustness of our transit timing analysis. Currently, all original ETD light curves are available on the VarAstro [^3] https://var.astro.cz/en/ server, which was launched last year.
3.3.2 ExoClock Project
The ExoClock project, initiated in September 2019, is an interactive platform designed to coordinate the regular monitoring of transiting exoplanets with small- to medium-sized telescopes. Its primary objective is to refine the ephemerides of targets for ESA’s upcoming Ariel mission, while also fostering international collaboration to improve the temporal coverage and precision of transit measurements.
@@ 3.3.2 ExoClock Project
3.3.3 Data Reduction of ETD and ExoClock
Before performing the light-curve analysis, we corrected for atmospheric effects inherent in ground-based observations. The out-of-transit (OOT) flux values were normalized close to unity by fitting a linear function of time to the OOT portion of each light curve. Following the procedure of I.-G. Jiang et al. (2016), we modeled the airmass effect with a third-degree polynomial and the seeing effect with a linear function. The original light curve, $F_{o}(t)$, was expressed as:
$F_{o}(t)=F(t)\mathcal{P}(t)\mathcal{Q}(s),$ (1)
where $F(t)$ is the corrected light curve, $\mathcal{P}(t)=a_{0}+a_{1}t+a_{2}t^{2}+a_{3}t^{3}$ represents the polynomial correction for airmass, and $\mathcal{Q}(s)=1+c_{0}s$ accounts for seeing, with s denoting the image-specific seeing value. Since seeing information was unavailable for the ETD data, we uniformly set $\mathcal{Q}(s)=1$. The polynomial coefficients ($a_{0},a_{1},a_{2},a_{3}$) were optimized via a numerical search to minimize the standard deviation, ensuring the OOT flux of $F(t)$ approached unity. This normalization was applied consistently to all light curves. In addition, we addressed intrinsic limitations of ETD data, such as the use of JD/UTC or HJD/UTC timestamps instead of BJD/TDB. To ensure accurate transit timing, all timestamps from the transits analyzed in this study were converted to BJD on the TDB scale, correcting for relativistic effects and Earth’s orbital motion (J. Eastman et al. 2010).
3.4 Published Ground-Based Light curves
As described above, in addition to the 7 new transit light curves and 250 light curves obtained from TESS, ETD, and the ExoClock Project, we have compiled and analyzed a total of 108 publicly available complete light curves. These include 3 from E. R. Adams et al. (2024), 2 from T. Chan et al. (2011), 21 from P. Leonardi et al. (2024), 2 from G. Maciejewski et al. (2011), 44 from G. Maciejewski et al. (2013), 28 from G. Maciejewski et al. (2016), and 8 from S. W. Yee et al. (2020). Additionally, we obtained light curves from the authors Hebb Lesbie, Efrain Alavarado, and Karen Collins via private communication and applied selection criteria to include only those that are complete, have an out-of-transit baseline, and possess timing uncertainties of less than 5 minutes. Finally, we incorporated 23 light curves from K. A. Collins et al. (2017), 2 from L. Hebb et al. (2009), and 1 from E. Alvarado et al. (2024), adding these 26 light curves to the existing database for analysis. The complete list of all 391 light curves is provided in Table 5.
@@ 3.4 Published Ground-Based Light curves
Object Number of light Sources Total number of
name curves taken light curves
@@ 34 Exoclock
44 G. Maciejewski et al. (2013)
28 G. Maciejewski et al. (2016)
23 K. A. Collins et al. (2017)
21 P. Leonardi et al. (2024)
@@ 3 UDRO
WASP-12  3 E. R. Adams et al. (2024) 391
 2 T. Chan et al. (2011)
 2 L. Hebb et al. (2009)
 2 G. Maciejewski et al. (2011)
 8 S. W. Yee et al. (2020)
@@ 1 DFOT
 1 E. Alvarado et al. (2024)
4 Transit Light-curve Fitting
For a precise and consistent transit-timing analysis, and to determine the individual mid-transit times ($T_{m}$) as well as to refine the stellar and planetary parameters of the WASP-12 system, we modeled all 391 available transit light curves using the Transit Analysis Package (TAP; J. Z. Gazak et al. 2012).
TAP employs the analytic transit model of K. Mandel & E. Agol (2002) and performs parameter estimation through a Markov Chain Monte Carlo (MCMC) framework, enabling robust determination of orbital and transit parameters. Apart from our newly acquired transits, the archival LCs were either collected directly from the respective publications or obtained by contacting the authors.
To ensure the homogeneity and reliability of the dataset, we applied the following selection criteria before incorporating any unpublished ground-based light curves, as well as published light curves from papers, ETD, and ExoClock, into our analysis:

### v2

@@ III.1.1 Data Reduction of high-precision TESS data
We retrieved TESS light curves from the NASA funded astronomical data archive repository Mikulski Archive for Space Telescopes (MAST) [^1] All the TESS data used in this paper can be found in MAST (MAST Team, 2021)., using the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) products generated by the Science Processing Operations Center (SPOC; Jenkins et al. 2016). Compared to standard SAP data, PDCSAP light curves reduce scatter, suppress short-term noise, and correct long-term instrumental trends (Smith et al. 2012; Stumpe et al. 2012, 2014; Caldwell et al. 2020).
@@
Transit windows were extracted within $\pm 0.125$ days of the predicted mid-transit times. Incomplete or noisy transits were excluded, as missing ingress or egress hampers precise timing (Barros et al. 2013). Data pre-processing, including extraction, normalization, and detrending, was performed with the JULIET package (Espinoza et al. 2019), which employs Bayesian inference through the MultiNest algorithm (Feroz et al. 2009, 2019; Buchner 2016). We retained only points with a quality flag of zero and converted time stamps to $\mathrm{BJD_{TDB}}$ by adding 2,457,000 (Eastman et al. 2010).
@@
Stellar and instrumental variability was modeled using a Gaussian Process (GP) with a Matérn kernel implemented in CELERITE (Espinoza et al. 2019). A normal prior was set for the mean out-of-transit flux, the dilution factor fixed to unity, and wide log-uniform priors adopted for GP hyperparameters (amplitude, timescale, and jitter). This procedure yielded GP-corrected light curves, optimized for precise transit timing and exoplanet characterization. Other details are given in Section IV.
III.2 New Ground-Based Light Curves
@@
For WASP-12 b, we obtained one transit on 7 January, 2025 using the 1.3 m Devasthal Fast Optical Telescope (DFOT) at the Devasthal, Nainital campus of ARIES. In addition, three transits were observed on 26 and 27 December, 2024 and 24 February, 2025 with the 0.61 m reflecting telescope, VASISTHA at the IERCOO campus of ICSP, Kolkata. Three further transits were recorded on 2 January, 4 January, and 26 March 2025 with the 0.3 m AG Optical IDK telescope at the Utah Desert Remote Observatory (UDRO) in Beryl, Utah. The exposure times were optimized to compensate for variable weather conditions. The data reduction procedures have been described in detail in Section III.2.4.
III.2.1 Devasthal Fast Optical Telescope
@@
The observations of one new transit of WASP-12 b were conducted using the 1.3-m Devasthal Fast Optical Telescope (DFOT), located at the Aryabhatta Research Institute of Observational Sciences (ARIES), Nainital, India. The DFOT, a modern Ritchey–Chrétien Cassegrain telescope with a 1.3 m aperture, was installed at Devasthal by DFM Engineering Inc., USA, and is operated by ARIES, an autonomous institute under the Department of Science and Technology (DST), Government of India. Equipped with a fast $f/4$ optical system, the telescope is mounted on a fork–equatorial system with single-axis tracking. Its secondary mirror is controlled by a five-axis actuator for precise focusing, while friction drives provide backlash-free motion in right ascension and declination, ensuring a pointing accuracy of $\sim 10$ arcsec (rms). For imaging, a back-illuminated CCD camera with $2048\times 2048$ pixels ($13.5\,\mu\mathrm{m}$ pixel size) and deep thermoelectric cooling down to $80^{\circ}$C was used. The transit observation using DFOT was carried out using an $R$-band filter.
III.2.2 Ionospheric and Earthquake Research Centre and Optical Observatory
@@
We observed three transits of the hot Jupiter WASP-12 b with the 0.61 m VASISTHA reflecting telescope at the Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), established by ICSP at Sitapur, Paschim Medinipur, West Bengal. This telescope, the largest in eastern India, is mounted on an Ascension 200 German equatorial system with high-resolution encoders. All observations were conducted in the Cousins $R$ band to reduce limb darkening, color-dependent extinction, and to allow high-cadence photometry (Holman et al. 2007). The telescope employs a 24 inch primary mirror with a focal ratio of $f/6.5$ and is equipped with an Atik 460EX Mono CCD camera. The detector has $2749\times 2199$ pixels of size $4.54~\mu\mathrm{m}$, providing an image scale of $0.235~\mathrm{arcsec~pixel}^{-1}$.
III.2.3 Utah Desert Remote Observatory
@@
III.2.4 Data Reduction of New Ground-based Data
@@
Before performing aperture photometry, we preprocessed the raw CCD images to remove instrumental and observational artifacts. We corrected defects such as bad pixels, cosmic ray hits, and pixel to pixel sensitivity variations to ensure reliable photometric measurements. Along with the science frames, calibration frames namely bias, dark, and flat fields are obtained to remove instrumental effects from the CCD images, including electronic offsets and thermal current, and to correct for pixel-to-pixel sensitivity variations and optical effects. Using IRAF [^2] The Image Reduction and Analysis Facility (IRAF) is a software package distributed by the National Optical Astronomy Observatory (NOAO), which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation (NSF). (Image Reduction and Analysis Facility), we apply standard routines such as zerocombine, flatcombine, and ccdproc under the imred package of NOAO to perform trimming, bias subtraction, dark correction, and flat–fielding, and we remove cosmic rays with the cosmic rays task in crutil. After calibration, we inspected the images for residual defects and alignment issues, displayed them with the task display, and interactively analyzed them using imexamine to estimate parameters such as the full width at half maximum (FWHM) of stellar point spread functions and the local sky background. These parameters are essential for defining the photometric aperture radius and sky annulus, which directly affect the precision of the extracted light curves in exoplanet transit analysis.
@@
: We performed the aperture photometry using the apphot and daophot packages in IRAF. First, we defined the photometry parameters in datapars, centerpars, fitskypars, and photpars to specify CCD characteristics, centering methods, sky background estimation, and aperture sizes. We prepared a coordinate file for the target star and nearby comparison stars, although stars could also be selected interactively. Aperture photometry was then applied across all science frames using the phot task to extract instrumental magnitudes, with magnitudes computed for the chosen aperture(s). The aperture radii were typically set to 2-3 times the full width at half maximum (FWHM) of the stellar point-spread function (PSF) and optimized to minimize scatter in the out-of-transit (OOT) light curve. We tracked stellar centroids across all frames, typically by cross-correlating with a reference image, to ensure precise aperture alignment during extraction. Photometry was then performed with the daophot package in IRAF on the target star and 2-8 nearby comparison stars of similar brightness and color. The aperture configuration and number of comparison stars were adjusted slightly from night to night according to atmospheric conditions and field orientation, following the approach of Jiang et al. (2016).
@@
Once we extracted the instrumental magnitudes, we compiled the results using txdump and related tools to obtain magnitudes, errors, and centroids for the target and comparison stars. For exoplanet transit studies, we performed differential photometry by constructing a reference flux from the summed fluxes of the chosen comparison stars and dividing the target star’s flux by this reference. This produced a time series of relative flux, forming the transit light curve. Then, we converted all time stamps to the Barycentric Julian Date in the Barycentric Dynamical Time system (BJDTDB) using the publicly available code of Eastman et al. (2010). Further modeling and fitting, allowed us to refine the stellar and planetary parameters of the exoplanet, which are described in detail in Section IV.
@@
Target Name Telescope Epoch TDB-based BJD Normalized Flux Normalized Flux Error
WASP-12 b 1.3 m 5652 2460684.12708776 1.0015969 0.0031616
WASP-12 b 1.3 m 5652 2460684.12749167 1.005571 0.0031616
WASP-12 b 1.3 m 5652 2460684.12789558 1.0047181 0.0031616
@@
III.3 Ground-based Observations from Public Databases
III.3.1 Exoplanet Transit Database
The ETD, established in September 2008, serves as a collaborative platform where amateur astronomers worldwide contribute transit measurements. As of January 4, 2026, the database contains 84,307 observations contributed by 1737 observers from different observatories across the globe. In addition to the TESS observations, we incorporated complete transit light curves from the Exoplanet Transit Database (ETD), selecting only those with a data quality index ($DQ<3$) to ensure consistency and reliability. These high-quality, community sourced datasets provided an essential complement to the TESS data, thereby improving the robustness of our transit timing analysis. Currently, all original ETD light curves are available on the VarAstro [^3] https://var.astro.cz/en/ server, which was launched last year.
III.3.2 ExoClock Project
The ExoClock project, initiated in September 2019, is an interactive platform designed to coordinate the regular monitoring of transiting exoplanets with small to medium-sized telescopes. Its primary objective is to refine the ephemerides of targets for ESA’s upcoming Ariel mission, while also fostering international collaboration to improve the temporal coverage and precision of transit measurements.
@@ III.3.2 ExoClock Project
III.4 Published Ground-Based Light curves
As described above, in addition to the 7 new transit light curves and 250 light curves obtained from TESS, ETD, and the ExoClock Project, we have compiled and analyzed a total of 108 publicly available complete light curves. These include 3 from Adams et al. (2024), 2 from Chan et al. (2011), 21 from Leonardi et al. (2024a), 2 from Maciejewski et al. (2011c), 44 from Maciejewski et al. (2013b), 28 from Maciejewski et al. (2016b), and 8 from Yee et al. (2020). Additionally, we obtained light curves from the authors Hebb Lesbie, Efrain Alavarado, and Karen Collins via private communication and applied selection criteria to include only those that are complete, have an out-of-transit baseline, and possess timing uncertainties of less than 5 minutes. Finally, we incorporated 23 light curves from Collins et al. (2017), 2 from Hebb et al. (2009), and 1 from Alvarado et al. (2024), adding these 26 light curves to the existing database for analysis. The complete list of all 391 light curves is provided in Table 5.
@@ III.4 Published Ground-Based Light curves
Number of light Sources
curves taken
@@ 34 Exoclock
44 Maciejewski et al. (2013b)
28 Maciejewski et al. (2016b)
23 Collins et al. (2017)
21 Leonardi et al. (2024a)
@@ 3 UDRO
 3 Adams et al. (2024)
 2 Chan et al. (2011)
 2 Hebb et al. (2009)
 2 Maciejewski et al. (2011c)
 8 Yee et al. (2020)
@@ 1 DFOT
 1 Alvarado et al. (2024)
IV Transit Light-curve Fitting
To perform a precise and homogeneous transit-timing analysis, and to determine the individual mid-transit times ($T_{\mathrm{m}}$) while refining the stellar and planetary parameters of the WASP-12 system, we modeled all 391 transit light curves using juliet (Espinoza et al., 2019). The transit light curves were modeled simultaneously with a Gaussian Process (GP) noise model. To do this, we adopted a simple, approximate Matérn kernel, implemented using the celerite formalism. The covariance between the $i$-th and $j$-th data points is described by
$k(\tau_{i,j})=\sigma_{GP}^{2}\,\tilde{M}(\tau_{i,j},\rho)+(\sigma_{i}^{2}+\sigma_{w}^{2})\,\delta_{i,j},$ (1)
where $k(\tau_{i,j})$ represents the $(i,j)$ element of the covariance matrix $\Sigma$, and $\tau_{i,j}=|t_{i}-t_{j}|$ is the temporal separation between the GP regressors $t_{i}$ and $t_{j}$, which in this work correspond to the observation times. The term $\sigma_{i}$ denotes the measurement uncertainty of the $i$-th data point, $\sigma_{GP}$ sets the GP amplitude (in ppm), and $\sigma_{w}$ (in ppm) represents an additional, unknown white-noise jitter component. The symbol $\delta_{i,j}$ denotes the Kronecker delta, which is equal to unity when $i=j$ and zero otherwise. The function $\tilde{M}(\tau_{i,j},\rho)$ is given by
$\tilde{M}(\tau_{i,j},\rho)=\left[(1+1/\epsilon)\exp\!\left(-[1-\epsilon]\sqrt{3}\tau/\rho\right)+(1-1/\epsilon)\exp\!\left(-[1+\epsilon]\sqrt{3}\tau/\rho\right)\right],$ (2)
which corresponds to the approximate Matérn component of the kernel and is characterised by a correlation length-scale $\rho$. This approach is particularly advantageous as it accounts for correlated (non-white) noise present during transit, thereby yielding more realistic uncertainties on the inferred planetary parameters. Within juliet, we simultaneously fitted the transit parameters and the GP hyperparameters, adopting appropriate priors and providing the relevant GP regressors. To correct for atmospheric transparency variations, we normalized the light curves by fitting a linear function to the out-of-transit (OOT) data during the fitting stage and including this OOT model simultaneously with the transit model. This approach avoids artificially removing degrees of freedom from the light-curve data and underestimating the uncertainties of the final parameters.
In addition to our newly obtained transit observations and the TESS, ETD, and ExoClock light curves, we used archival light curves obtained either from the published literature or through private communication with the respective authors.
To ensure the homogeneity and reliability of the dataset, we applied the following selection criteria before incorporating any unpublished ground-based light curves, as well as published light curves from the literature, ETD, and ExoClock, into our analysis: